# Naming, Describing, and Quantifying Visual Objects in Humans and LLMs

### v1

@@ Abstract
While human speakers use a variety of different expressions when describing the same object in an image, giving rise to a distribution of plausible labels driven by pragmatic constraints, the extent to which current Vision & Language Large Language Models (VLLMs) can mimic this crucial feature of language use is an open question. This applies to common, everyday objects, but it is particularly interesting for uncommon or novel objects for which a category label may be lacking or fuzzy. Furthermore, humans show clear production preferences for highly context-sensitive expressions, such as the quantifiers ‘few’ or ‘most’. In our work, we evaluate VLLMs (FROMAGe, BLIP-2, LLaVA) on three categories (nouns, attributes, and quantifiers) where humans show great subjective variability concerning the distribution over plausible labels, using datasets and resources mostly under-explored in previous work. Our results reveal mixed evidence on the ability of VLLMs to capture human naming preferences, with all models failing in tasks that require high-level reasoning such as assigning quantifiers.
@@ 1 Introduction
Figure 1: Datasets used and distribution of human answers/labels. In NOUN (left), we focus on the frequency of color and texture attributes in the descriptions. In ManyNames (middle), each object is associated with the frequency of the nouns used to describe it. In QUANT (right), each image is associated with a probability distribution over a list of quantifiers that humans selected when answering ‘How many of the objects are animals?’.
@@
In our work, we investigate this issue in three different production conditions. First of all, we consider the ManyNames dataset (Silberer et al., 2020a, b), where annotators assign labels to describe common objects in images in a referential expression generation setting (Yu et al., 2016; Kazemzadeh et al., 2014). We also explore two additional resources that have not received much attention within the NLP community and that allow us to broaden the horizons of this phenomenon. First, we analyze the NOUN dataset (Horst and Hout, 2016), where speakers describe uncommon and novel objects: we focus on both the choice of the adjectives and how they distribute in the across-subject distribution. Finally, we investigate human production variability arising from the context-sensitive nature of non-numerical quantifiers using the data collected by Pezzelle et al. (2018).
We evaluate three VLLMs (FROMAGe, BLIP-2, LLaVA) on the above-mentioned tasks in a zero-shot setting. We sample multiple times from the model using nucleus sampling, mimicking various human speakers, and compare the generated samples against human production patterns using different metrics (Jensen–Shannon divergence and Pearson’s correlation, depending on the task at hand). Our results show that models weakly to moderately mimic human distributions in naming common and uncommon objects. Instead, they fail to mimic human distributions when selecting quantifiers, as highlighted by our in-depth analyses.
@@ 2 Tasks and Datasets
The Novel Object and Unusual Name (NOUN) dataset (Horst and Hout, 2016) contains 64 images of multipart, multicolored, and three-dimensional uncommon and novel objects. The dataset was originally created for behavioral studies on word learning and, to the best of our knowledge, it has not been used for NLP research. We focus on the naming task, where participants were asked to answer the question “What would you call this object?”. The answers are sentences like: ‘a plastic object with red stuff on top’. For each object, the proportion of colors (e.g., ‘red’, ‘bronze’) and textures (e.g., ‘soft’, ‘rough’) was calculated as the number of attributes given the number of responses. An example from the dataset is reported in Figure 1 (left), together with the ratio of colors and textures in human responses. In NOUN, we examine human production preferences on a high level, by just looking at the frequency according to which certain adjectives are used.
@@
In ManyNames (Silberer et al., 2020a, b), the authors collected names for 25K objects from VisualGenome (Krishna et al., 2017) by asking human annotators. Each object is associated with an average number of 35.3 annotations. More than 90% of the objects are associated with more than one unique label (5.7 average name types per object). An example is shown in Figure 1 (middle). When describing the object in the red box, most annotators referred to it as ‘chair’, while around 30% said ‘sofa’, and the remaining ones used ‘couch’ and ‘armchair’. The images in ManyNames are classified into 7 domains (e.g., vehicles, people, etc.): for computational constraints, we evaluated 300 randomly sampled objects from each domain. Differently from NOUN, here we examine production preferences on a more fine-grained level using the actual distribution over multiple labels.
@@
To study how quantifiers are used when referring to quantities grounded in images, Pezzelle et al. (2018) introduced a dataset of visual abstract scenes containing a variable number of animals and artifacts and asked human participants to answer the question “How many of the objects are animals?". Participants could select the answer from a list of nine quantifiers: ‘none’, ‘almost none’, ‘the smaller part’, ‘few’, ‘some’, ‘many’, ‘most’, ‘almost all’, ‘all’. The authors used images with 17 different proportions of animals and artifacts (ranging from 0% to 100%). In our work, we tested 50 images for each of the 17 proportions in the dataset, resulting in a total number of 850 images. [^1] The actual images used in our experiment come from Testoni et al. (2019), which built a large-scale dataset using the stimuli and pipeline by Pezzelle et al. (2018).
@@ 3.1 Generation
In our work, we test the performance of three models in a zero-shot setting: BLIP-2 (Li et al., 2023b), FROMAGe (Koh et al., 2023), and LLaVA (Liu et al., 2023b, a). All three models can be prompted for zero-shot generation. Additional details are discussed in Appendix A.4. For each of the three tasks described in Section 2, we used prompts that resembled the instructions provided to human annotators during the dataset collection. ManyNames: Q: Please name the object in the red box with the first name that comes to mind. A:. NOUN: Q: What do you see in the image? A:. QUANT: Question: How many animals are there in the image? Pick the best among the following options: , followed by the list of the nine quantifiers, each associated with a letter (from A to I). The ordering of the quantifiers is randomized at each inference step. We sample multiple times from each model using nucleus sampling decoding (Holtzman et al., 2019), with $p=0.9$, $t=0.5$ (different hyperparameter configurations did not significantly affect the overall results). For each task, we sample the model 20 times and filter out ill-formed answers, such as empty strings or question repetitions. After filtering, we randomly take 10 generations per image for ManyNames and NOUN, and 15 for QUANT. In this way, we have the same number of generations for each image/object. We will release our code upon acceptance at: github.com/anonymous/repo.
@@ 3.2 Evaluation
Each object in NOUN is associated with color and texture saliency, i.e., how often speakers described the object using these attributes. We use a string-match approach (see Appendix A.2) to analyze the model output and compute color and texture saliency. We compute the Pearson’s r correlation between human and model saliency, considering all objects. Each object in ManyNames is associated with $H$ unique nouns assigned by human annotators and $M$ unique nouns sampled from the model output, together with their frequency. Given $A=H\cup M$, we construct two term-frequency vectors for human and model output, $h$ and $m$, respectively, with $|h|=|m|=|A|$. Each noun in $A$ is mapped to a unique position in $h$ and $m$ and each vector is filled with its normalized frequency. We evaluate the models by computing the inverse Jensen–Shannon (JS, bounded between 0 and 1) divergence (Lin, 1991) between $h$ and $m$. See Figure 5 in the Appendix for an example. Each image in QUANT is associated with a probability distribution over 9 quantifiers, depending on the proportion of animals and artifacts. From the model outputs, we extract the relative frequency of each quantifier and compute Pearson’s r correlation with the human distribution. We then average the correlation results over all images. Correlation is bounded between -1 and 1. Higher is better for all the metrics.
Figure 2: NOUN, QUANT: correlation between human responses and model samples. ManyNames: inverse JS divergence between the frequency of the nouns chosen by annotators and the ones generated by the model.
@@ 3.3 Results
As we can observe from Figure 2, the results for ManyNames and NOUN (color saliency) show a clear trend: all the models correlate, to some extent, with human production, with LLAVA obtaining the highest correlations for both tasks (around 0.5) and significantly outperforming (t-test, $p<0.01$) both BLIP2 and FROMAGe. [^2] Appendix A.1 shows per-domain results for ManyNames. These findings combine with previous work showing the primacy of LLaVA over other models (Liu et al., 2023b, a) However, the remaining tasks show critical weaknesses for all models. First, none of the models achieve a statistically significant correlation for texture saliency ($p>0.05$). We conjecture that texture attributes are less common for the models compared to colors, and thus they may be less accurate when generating them: we leave an in-depth analysis of this issue for future work. Despite the correlation results being similar across models, our manual inspection reveals interesting differences: while the low performance of FROMAGe is due to an under-generation of texture attributes, the opposite is true for LLaVA, with BLIP-2 being more flexible in terms of texture attribute generation but not aligned with human variability (see Figure 6 for an example). Finally, all models show almost no correlation in assigning quantifiers to visual scenes, highlighting a severe limitation of all models on this task. We scrutinize this issue in the following Section.
Figure 3: Correlation results (y-axis) broken down by proportion of targets (animals) in the image (x-axis).
@@ 4 The Curious Case of Quantifiers
We run some analyses to investigate the poor performance of all models in the QUANT task. First of all, we acknowledge that the multiple-choice prompting used in QUANT is different and more complex than the prompts used for the other datasets. Still, it is unlikely that this is the main reason behind the poor performance of all models. Figure 3 shows the correlation results broken down by the proportion of animals in the image. We observe that even though the overall correlation results are similar across models (Figure 2), models perform quite differently depending on the proportion of animals in the scene. While BLIP-2 performs relatively well on the ‘extreme’ proportions (no animals or all animals in the image, when speakers generally choose the quantifiers ‘none’ and ‘all’, respectively), LLaVA excels at intermediate proportions, and FROMAGe performs better on proportions above 50%. Can we conclude that models properly perform the task of assigning the most likely quantifiers for some proportions? These results evoke two hypotheses: (a) The models are capable of selecting plausible quantifiers only for some proportions or, vice versa, they understand only some of the quantifiers analyzed; (b) The models have a bias towards some specific quantifiers, regardless of the proportion of targets in the scene, leading to a decent perform on some proportions as a side effect. Our additional analyses, reported in Figures 7 and 8 in the Appendix, support hypothesis (b): FROMAGe has a strong bias towards selecting the quantifier ‘many’; BLIP-2 frequently selects the extreme quantifiers ‘none’ and ‘all’, and its selection is not influenced by the proportion of targets; LLaVA has a bias towards selecting the quantifier ‘some’, regardless of the proportion of targets. To further shed light on this result, we qualitatively assess the ‘counting’ skills of the models, a crucial skill to succeed in assigning quantifiers. As the examples in Figure 9 in the Appendix illustrate, all models struggle to successfully count how many animals appear in the image. We hypothesize that the reason for the poor performance in assigning quantifiers lies in the quantity estimation and comparison skills of the models.
@@ 5 Conclusion
In the following, we discuss some limitations of our study that may inspire follow-up work in this direction. The poor performance on the quantification tasks may stem from the higher complexity of the prompt used (multiple choice prompting). Even though in our paper we discuss how analyzing the output variability allows us to gain valuable insights even when the model is not accurate, we can not rule out the possibility that a simpler prompt may lead to more accurate results. As an initial step, we used a prompt that corresponds to the instruction provided to the participants of the original experiment in Pezzelle et al. (2018).
@@ A.3 QUANT Appendix
Figures 7 and 8 show additional analyses on the QUANT dataset. They are discussed in Section 4. Figure 9 shows some qualitative examples of the models’ output when asked to answer the question ‘How many animals are there in the image?’. Note that the images are randomly selected from QUANT. QUANT is released without a specific license.
@@ A.4 Models Appendix
While FROMAGe is trained with a contrastive learning objective for image captioning and it is shown to perform particularly well with longer textual contexts, BLIP-2 jointly optimizes three pre-training objectives that share the same input format and model parameters: image-text contrastive learning, image-grounded text generation, and image-text matching. The main innovation of LLaVA is the use GPT-4 generated visual instruction tuning data. Moreover, LLaVA has a simpler scheme to connect image and language representations compared to BLIP-2 and FROMAGe. We used blip2-opt-2.7b and llava-v1.5-7b, while for FROMAGe we used the model made available by Koh et al. (2023). FROMAGe and LLaVA are released with Apache-2.0 license. BLIP-2 is distributed with BSD 3-Clause License. We run our experiments in accordance with the model license.
A.5 Additional Details

### v3

@@ Abstract
While human speakers use a variety of different expressions when describing the same object in an image, giving rise to a distribution of plausible labels driven by pragmatic constraints, the extent to which current Vision & Language Large Language Models (VLLMs) can mimic this crucial feature of language use is an open question. This applies to common, everyday objects, but it is particularly interesting for uncommon or novel objects for which a category label may be lacking or fuzzy. Furthermore, similar patterns of variation are observed among human speakers for highly context-sensitive expressions, such as the quantifiers ‘few’ or ‘most’. In our work, we evaluate VLLMs (FROMAGe, BLIP-2, LLaVA) on three categories (nouns, attributes, and quantifiers) where humans show great subjective variability concerning the distribution over plausible labels, using datasets and resources mostly under-explored in previous work. Our results reveal mixed evidence on the ability of VLLMs to capture human naming preferences at generation time: while some models are good at mimicking human distributions for nouns and attributes, all of them fail to assign quantifiers, a task that requires more accurate, high-level reasoning.
@@ 1 Introduction
Figure 1: Datasets used in our experiments and distribution of human answers/labels. In NOUN (left), we focus on the frequency of color and texture attributes in the generated descriptions. In ManyNames (middle), each object is associated with the frequency of the nouns used to describe it. In QUANT (right), each image is associated with a probability distribution over a list of quantifiers that humans selected when answering the question ‘How many of the objects are animals?’.
@@
In our work, we investigate this issue, which has recently entered the NLP research community (Plank, 2022), in three different production conditions. First of all, we consider the ManyNames dataset (Silberer et al., 2020a, b), where annotators assign labels to describe common objects in images in a referential expression generation setting (Yu et al., 2016; Kazemzadeh et al., 2014). We also explore two additional resources that have not received much attention within the NLP community and that allow us to broaden the horizons of this phenomenon. First, we analyze the NOUN dataset (Horst and Hout, 2016), where speakers describe uncommon and novel objects: we focus on both the choice of the adjectives and how they distribute in the across-subject distribution. Finally, we investigate human production variability arising from the context-sensitive nature of non-numerical quantifiers using the data collected by Pezzelle et al. (2018).
We evaluate three VLLMs (FROMAGe, BLIP-2, LLaVA) on the above-mentioned tasks in a zero-shot setting. We sample multiple times from the model using nucleus sampling, mimicking various human speakers, and compare the generated samples against human production patterns using different metrics (Jensen–Shannon divergence and Pearson’s correlation, depending on the task at hand). Our results show that models weakly to moderately mimic human distributions in naming common and uncommon objects. Instead, all of them fail to mimic human distributions when selecting quantifiers, as highlighted by our in-depth analyses.
@@ 2 Tasks and Datasets
The Novel Object and Unusual Name (NOUN) dataset (Horst and Hout, 2016) contains 64 images of multipart, multicolored, and three-dimensional uncommon and novel objects. The dataset was originally created for behavioral studies on word learning and, to the best of our knowledge, it has not been used for NLP research. We focus on the naming task, where participants were asked to answer the question “What would you call this object?”. The answers are sentences like: ‘a plastic object with red stuff on top’. For each object, the proportion of colors (e.g., ‘red’, ‘bronze’) and textures (e.g., ‘soft’, ‘rough’) was calculated as the number of attributes given the number of responses. An example from the dataset is reported in Figure 1 (left), together with the ratio of colors and textures in human responses. In NOUN, we examine human production preferences on a high level, by looking at the frequency according to which certain adjectives (related to color and texture attributes) are used.
Figure 2: Examples of the output generated by LLaVA (multiple samples with nucleus sampling decoding) for the three tasks analyzed in our work. For each task, a sample of the answers provided by the model is displayed. For space constraints, we only report a few random samples for each task.
@@
In ManyNames (Silberer et al., 2020a, b), the authors collected names for 25K objects appearing in real-world images from VisualGenome (Krishna et al., 2017) by asking human annotators to generate a name for them. Each object (highlighted by a red box in the image) is associated with an average number of 35.3 annotations. More than 90% of the objects are associated with more than one unique label (5.7 average name types per object). An example is shown in Figure 1 (middle). When describing the object in the red box, most annotators referred to it as ‘chair’, while around 30% said ‘sofa’, and the remaining ones used ‘couch’ and ‘armchair’. The images in ManyNames are classified into 7 domains (e.g., vehicles, people, animals, etc.): for computational constraints, we evaluated 300 randomly sampled objects from each domain. Different from NOUN, we examine production preferences on a more fine-grained level using the actual distribution over multiple labels.
@@
To study how quantifiers are used when referring to quantities grounded in images, Pezzelle et al. (2018) introduced a dataset of visual abstract scenes containing a variable number of animals and artifacts and asked human participants to answer the question “How many of the objects are animals?". Participants could select the answer from a list of nine pre-selected quantifiers: ‘none’, ‘almost none’, ‘the smaller part’, ‘few’, ‘some’, ‘many’, ‘most’, ‘almost all’, and ‘all’. The authors used images with 17 different proportions of animals and artifacts (ranging from 0% to 100%). In our work, we tested 50 images for each of the 17 proportions in the dataset, resulting in a total number of 850 images. [^1] The actual images used in our experiment come from Testoni et al. (2019), which built a large-scale dataset using the stimuli and pipeline by Pezzelle et al. (2018).
@@ 3.1 Generation
In our work, we test the performance of three models in a zero-shot setting: BLIP-2 (Li et al., 2023b), FROMAGe (Koh et al., 2023), and LLaVA 1.5 (Liu et al., 2023b, a). All three models can be prompted for zero-shot generation. Additional details are discussed in Appendix A.4. For each of the three tasks described in Section 2, we used prompts that resembled the instructions provided to human annotators during the dataset collection. ManyNames: Q: Please name the object in the red box with the first name that comes to mind. A:. NOUN: Q: What do you see in the image? A:. QUANT: Question: How many animals are there in the image? Pick the best among the following options: , followed by the list of the nine quantifiers, each associated with a letter (from A to I). The ordering of the quantifiers is randomized at each inference step. Although investigating several variations of the above-mentioned prompts is beyond the scope of the paper, we discuss some insights on this aspect in Appendix A.5. We sample multiple times from each model using nucleus sampling decoding (Holtzman et al., 2019), with $p=0.9$, $t=0.5$ (different hyperparameter configurations did not significantly affect the overall results, as discussed in Appendix A.5). For each task, we sample the model 20 times and filter out ill-formed answers, such as empty strings or question repetitions. After filtering, we randomly take 10 generations per image for ManyNames and NOUN, and 15 for QUANT. In this way, we have the same number of generations for each image/object. Some examples of the output generated by LLaVa for the three tasks analyzed are reported in Figure 2. We release our code at: https://github.com/albertotestoni/ndq_visual_objects.
Figure 3: For NOUN and QUANT, the plot shows the correlation between human responses and model samples. For ManyNames, it shows the inverse JS divergence between the frequency of the nouns chosen by annotators and the ones generated by the model.
@@ 3.2 Evaluation
Each object in NOUN is associated with color and texture saliency, i.e., how often speakers described the object using these attributes. We use a string-match approach (see Appendix A.2) to analyze the model output and compute color and texture saliency. We then compute the Pearson’s r correlation between human and model saliency, considering all objects.
Each object in ManyNames is associated with $H$ unique nouns assigned by human annotators and $M$ unique nouns sampled from the model output, together with their frequency. Given $A=H\cup M$, we construct two term-frequency vectors for human and model output, $h$ and $m$, respectively, with $|h|=|m|=|A|$. Each noun in $A$ is mapped to a unique position in $h$ and $m$ and each vector is filled with its normalized frequency. We evaluate the models by computing the inverse Jensen–Shannon (JS, bounded between 0 and 1) divergence (Lin, 1991) between $h$ and $m$. See Figure 6 in the Appendix for an example.
Each image in QUANT is associated with a probability distribution over 9 quantifiers, depending on the proportion of animals and artifacts. From the model outputs, we extract the relative frequency of each quantifier and compute Pearson’s r correlation with the human distribution. We then average the correlation results over all images. Correlation is bounded between -1 and 1. Higher is better for all the metrics.
@@ 3.3 Results
As we can observe from Figure 3, the results for ManyNames and NOUN (color saliency) show a clear trend: all the models correlate, to some extent, with human production, with LLAVA obtaining the highest correlations for both tasks (around 0.5) and significantly outperforming (t-test, $p<0.01$) both BLIP2 and FROMAGe. [^2] Appendix A.1 shows per-domain results for ManyNames. These findings align with previous work showing the primacy of LLaVA over other models (Liu et al., 2023b, a). However, the remaining tasks show critical weaknesses for all models. First, none of the models achieve a statistically significant correlation for texture saliency (all have $p>0.05$). We conjecture that texture attributes are less common for the models compared to colors, and thus they may be less accurate when generating them: we leave an in-depth analysis of this issue for future work. Despite the correlation results being similar across models, our manual inspection reveals interesting differences: while the low performance of FROMAGe is due to an under-generation of texture attributes, the opposite is true for LLaVA, with BLIP-2 being more flexible in terms of texture attribute generation but not aligned with human variability (see Figure 7 for an example). Finally, all models show almost no correlation in assigning quantifiers to visual scenes, highlighting a severe limitation of all models on this task. We scrutinize this issue in the following Section.
Figure 4: Pearson correlation results (y-axis) broken down by proportion of targets (animals) in the image (x-axis) in the QUANT task and dataset.
@@ 4 The Curious Case of Quantifiers
We run some analyses to investigate the poor performance of all models in the QUANT task. First of all, we acknowledge that the multiple-choice prompting used in QUANT is different and more complex than the prompts used for the other datasets. Still, it is unlikely that this is the main reason behind the poor performance of all models.
Figure 4 shows the correlation results broken down by the proportion of animals in the image. We observe that even though the overall correlation results are similar across models (Figure 3), they perform quite differently depending on the proportion of animals in the scene. While BLIP-2 performs relatively well on the ‘extreme’ proportions (no animals or all animals in the image, when speakers generally choose the quantifiers ‘none’ and ‘all’, respectively), LLaVA excels at intermediate proportions, and FROMAGe performs better on proportions above 50%. Can we conclude that models properly handle the task of assigning the most likely quantifiers for some proportions? These results evoke two hypotheses: (a) The models are capable of selecting plausible quantifiers only for some proportions or, vice versa, they understand only some of the quantifiers analyzed; (b) The models have a bias towards some specific quantifiers, regardless of the proportion of targets in the scene, leading to a decent perform on some proportions as a side effect. Our additional analyses, reported in Figures 8 and 9 in the Appendix, support hypothesis (b): FROMAGe has a strong bias towards selecting the quantifier ‘many’; BLIP-2 frequently selects the extreme quantifiers ‘none’ and ‘all’, and its selection is not influenced by the proportion of targets; LLaVA has a bias towards selecting the quantifier ‘some’, regardless of the proportion of targets.
To further shed light on this result, we qualitatively assess the ‘counting’ skills of the models, a crucial skill to succeed in assigning quantifiers. As the examples in Figure 11 in the Appendix illustrate, all models struggle to successfully count how many animals appear in the image. We hypothesize that the reason for the poor performance in assigning quantifiers lies in the quantity estimation and comparison skills of the models. This observation is in line with recent research investigating the poor ‘counting’ skills of current models (Paiss et al., 2023).
@@ 5 Conclusion
In the following, we discuss some limitations of our study that may inspire follow-up work in this direction. The poor performance on the quantification tasks may stem from the higher complexity of the prompt used (multiple choice prompting). Even though in our paper we discuss how analyzing the output variability allows us to gain valuable insights even when the model is not accurate, we can not rule out the possibility that a simpler prompt may lead to more accurate results. As an initial step, we used a prompt that corresponds to the instruction provided to the participants of the original experiment in Pezzelle et al. (2018). In Appendix A.5 we discuss the effect of re-phrasing the original prompt instructions.
@@
Acknowledgments
We would like to thank the Dialogue Modelling Group (DMG) at the University of Amsterdam for their feedback and support at the different stages of this work. We thank Brent Brakenhoff for his input on some preliminary experiments with the ManyNames dataset. Alberto Testoni is supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 819455, PI R. Fernández).
@@ A.3 QUANT Appendix
Figures 8 and 9 show additional analyses on the QUANT dataset. They are discussed in Section 4. Figure 11 shows some qualitative examples of the models’ output when asked to answer the question ‘How many animals are there in the image?’. The images are randomly selected from QUANT. QUANT is released without a specific license.
@@
Figure 10: The role of different top_p values for nucleus sampling decoding using the LLaVa model.
@@ A.4 Models Appendix
While FROMAGe is trained with a contrastive learning objective for image captioning and it is shown to perform particularly well with longer textual contexts, BLIP-2 jointly optimizes three pre-training objectives that share the same input format and model parameters: image-text contrastive learning, image-grounded text generation, and image-text matching. The main innovation of LLaVA is the use GPT-4 generated visual instruction tuning data. Moreover, LLaVA has a simpler scheme to connect image and language representations compared to BLIP-2 and FROMAGe. We used blip2-opt-2.7b and llava-v1.5-7b, while for FROMAGe we used the model made available by Koh et al. (2023). FROMAGe and LLaVA are released with an Apache-2.0 license. BLIP-2 is distributed with BSD 3-Clause License. We run our experiments under the model license.
A.5 Generation Details
The Effect of Different top_p Values
We experimented with various top_p values for nucleus sampling decoding. As illustrated in Figure 10 (showing the results for LLaVA, with similar results for the other models), we observe that this variable does not play a significant role in our experimental setup for all the tasks analyzed.
Different prompts
In our experiments, we prompted the models with the same instructions provided to human annotators during the collection of the different datasets analyzed. We also experimented with small variations of the above-mentioned prompts, such as ‘What is the object in the image?’ for the NOUN dataset, ‘Name the object in the red box with the most appropriate single name’ for ManyNames, and a more detailed instruction for QUANT, such as ‘Carefully examine the image. Can you determine the proportion of animals present, compared to objects? Please select the most accurate answer from the options below’. While we do not observe any significant difference between NOUN and ManyNames, the revised prompt for QUANT leads to a slight improvement in the model performance, with LLaVA reaching a correlation of 0.29. Still, the low absolute correlation coefficient highlights that computational models struggle to accurately assign quantifiers to visual scenes. This result demonstrates that the prompt may influence the performance of the models on this task. Although exploring which prompts work best was beyond the scope of this paper, we leave a systematic exploration of this aspect to future research.
A.6 Additional Details